Q1. Firm XYZ is currently financed entirely with equity. The market value of the firm's assets and equity is ?? = ?? = 500, and the expected return on the firm's assets and equity is ?? = ?? = 12.5 percent. Suppose the firm issues debt with a value of ? = 200, and uses the proceeds to retire equity. The market value of the firm remains the same, ?? = ?? + ? = 500. If the expected return on the debt is ?? = 7 percent, what is the expected return on the firm's levered equity?
Answer #1

Debt Equity ratio = 200/ (500-200) = 2/3

expected return on the firm's levered equity = Ru + (D/E) * (Ru-Rd)

= 12.5%+ (2/3)*(12.5%- 7%)

=16.17%